# Critical Learning Periods: Leveraging Early Training Dynamics for Efficient Data Pruning

### v1

@@ Abstract
Neural Machine Translation models are extremely data and compute-hungry. However, not all data points contribute equally to model training and generalization. Data pruning to remove the low-value data points has the benefit of drastically reducing the compute budget without significant drop in model performance. In this paper, we propose a new data pruning technique: Checkpoints Across Time (CAT), that leverages early model training dynamics to identify the most relevant data points for model performance. We benchmark CAT against several data pruning techniques including COMET-QE, LASER and LaBSE. We find that CAT outperforms the benchmarks on Indo-European languages on multiple test sets. When applied to English-German, English-French and English-Swahili translation tasks, CAT achieves comparable performance to using the full dataset, while pruning up to 50% of training data. We inspect the data points that CAT selects and find that it tends to favour longer sentences and sentences with unique or rare words.
@@ 1 Introduction
In recent years, there has been significant improvement in the quality of Neural Machine Translation (NMT) models (Johnson et al., 2017; Arivazhagan et al., 2019; Liu et al., 2020; Fan et al., 2021; Reid et al., 2021; Ramesh et al., 2022; Costa-jussà et al., 2022; Bapna et al., 2022; Gala et al., 2023). The success of these models is primarily due to many factors including pretraining data size, compute abundance and the ever-increasing model size. Despite the gains due to growing datasets, large dataset sizes have also posed significant hurdles for maintaining data quality. Significant portions of web-scraped data used for language model pretraining have been shown to be of low quality, machine-generated spam, pornographic content (Kreutzer et al., 2022). Given the ever-growing size of parallel corpora, it often becomes laborious for humans to assess the quality at scale (Freitag et al., 2021; Agarwal et al., 2022c; Longpre et al., 2023). Quality of pretraining data for large NMT models often face amplified quality issues because of extensive bitext mining approaches to automatically extracting translation pairs from either monolingual corpora (Guo et al., 2018; Schwenk et al., 2021a, b; Costa-jussà et al., 2022; Ramesh et al., 2022; Vegi et al., 2022; Gala et al., 2023) or document-aligned corpora (Bañón et al., 2020; Steingrímsson et al., 2021; Ramesh et al., 2022; Gala et al., 2023). Bitext mining approaches, although far more cost-effective than human curation, have been shown to introduce significant noise into datasets by relying on sub-optimal sentence embedding models (Thompson & Koehn, 2019; Feng et al., 2022; Heffernan et al., 2022) to match sub-optimal translation pairs (Kreutzer et al., 2022).
@@
CAT is motivated by fundamental work in machine learning which has shown that there are distinctive periods to learning in deep neural network training. Work shows that more common, easier features are learnt earlier in training, with the most challenging features learnt in the last stages of training (Jiang et al., 2020; Achille et al., 2017; Siddiqui et al., 2022). Prior work has also shown that variance in gradients with respect to inputs (Agarwal et al., 2022a) or gradient patterns themselves vary depending on the period of learning (Faghri et al., 2020). However, gradients remain expensive to leverage as a learning signal to identify relevant instances. Furthermore, almost the entirety of this work has focused on a computer vision setting, with the exception of Swayamdipta et al. (2020) which uses differences in the model’s confidence in the true class, and the variability of this confidence across epoch to cluster data. Here – we explore applying the insights that there are distinctive periods of training to an NLP setting and leveraging for the purpose of pruning. We propose a more efficient method to compute signal relative to prior methods which have required computing gradients instead. Our hypothesis is that those with the largest variability indicate the easy examples which show the quickest learning and gains in certainty early in training.
@@
We study the effectiveness of CAT across diverse linguistic contexts. Specifically, we utilise English-German and English-French pairs from the Indo-European family using WMT datasets, and in the English-Swahili pair from the Bantu family with automatically aligned datasets. Through a series of extensive experiments, we show that CAT results in significant improvements in translation quality over random pruning (randomly selecting $X\%$ subset of the training data) and other embedding-based quality estimation methods such as LaBSE (Feng et al., 2022). We summarize our key contributions below:
@@
    We propose a novel method, Checkpoints Across Time (CAT), that leverages perplexity variation to prune of MT datasets with significant improvements over existing methods for Indo-European languages. We show that CAT results in significant improvements in translation quality over random pruning (randomly selecting $X\%$ subset of the training data) and other embedding-based quality estimation methods such as LaBSE (Feng et al., 2022). Using our best CAT-based method on English-German, English-French and English-Swahili translation tasks, we are able to achieve comparable performance while pruning up to 50% of the training data.
@@ 2.1 Data Pruning in ML
Data pruning has been extensively studied in the work of (Marion et al., 2023; Li et al., 2018; Raju et al., 2021; Swayamdipta et al., 2020; Boubdir et al., 2023). Li et al. (2018); Sorscher et al. (2022) and Raju et al. (2021) focus on data pruning for computer vision whereas Swayamdipta et al. (2020) focuses on data pruning for language tasks. Various techniques have been proposed for data pruning, such as utilization of loss profile (Siddiqui et al., 2022), gradients confidence score (Agarwal et al., 2022c), and data classification using a model (Swayamdipta et al., 2020). Increasingly techniques have been extended from primarily computer vision settings (Sorscher et al., 2023) to more recently expanding the treatment of data pruning language settings (Marion et al., 2023). These methods quantify the importance of data points to model generalization. Our work follows along these lines of identifying critical data points by leveraging the difference in perplexity between early checkpoints.
@@ 2.2 Model-based Data Pruning for NMT
All these approaches rely on the existence of a small trusted data set or trusted models to extrinsically define what “good data” looks like. Our approach, while also using model perplexity as a metric, however, does not require such pre-selection, and focuses on model training dynamics instead. Thereby, it can be applied more broadly. Our experiments aim at this less constrained setting without any given high-quality data, so we compare against model-based selection with a MT-QE model and a multilingual LLM.
@@ 2.4 Comparison Studies
With this plethora of filtering methods and application scenarios, there have been few works to systematically compare them. The WMT shared tasks for corpus filtering (Koehn et al., 2018, 2019, 2020) have shed a light on how different languages and resourcedness conditions pose different challenges for above described filtering methods in practice, and how they affect downstream NMT performance. Bane & Zaretskaya (2021) compare data filtering techniques with a more qualitative angle: They find that NMT scores and MUSE embeddings have the highest correlation with human quality judgments, NMT scores works best for in-domain selection, and MUSE/XLM-R for out-of-domain generalization. Herold et al. (2022) compare multiple data noise detection techniques according to their effectiveness for filtering out specific types of data noise. They find that LASER does not detect incorrect languages, and both LASER and cross-entropy perform weakly on detecting misaligned sentences and over/undertranslation. Similarly, Bane et al. (2022) find cross-entropy filtering empirically superior to other methods (XLM-R, MUSE, LASER, COMET) for filtering out various synthetic noise types (Khayrallah & Koehn, 2018). In their setup, COMET fails at filtering out misaligned sentences, but shows particular sensitivity for target-side omissions, NMT scoring for source-side omissions, while LASER and COMET do not filter mismatching numbers well. Dakwale et al. (2022) compare LASER and LaBSE on low-resource corpus filtering tasks, and find them competitive if languages are included in the pretraining of the underlying embedding models.
@@ 3 CAT: Checkpoints Across Time
Our goal is to extract the most valuable data points that would contribute to model generalization and train the model solely on this. Our approach, while simple, leverages the difference in learning stages in deep neural network optimization. Firstly, we train a NMT model on the full dataset for only a few epochs, ideally a small fraction of the total time needed for the model to converge. Subsequently, we use the checkpoints from this initial stage to compute the perplexity of each data point. These perplexity scores serve as the basis for selecting candidate data points for subsequent pruning using different perplexity profiles. This approach draws upon the findings of Swayamdipta et al. (2020) that categorize training data relevance based on the learning dynamics during training, and Agarwal et al. (2022b) which leverage the difference in gradients across training to identify easy versus challenging examples. Unlike Agarwal et al. (2022b) which uses computationally expensive variance of gradients, we use perplexity. For CAT we can utilize the perplexity scores in two ways:
@@ 3.1 CAT-DIFF
From the ranked data points, we select those with the highest differences between early epochs  [^1] Either 5th and 1st or 2nd and 1st., according to a threshold at a chosen percentile. For example, for pruning of 50% of data points, we keep all data points that rank higher than the 50th percentile.
@@ 3.2 CAT-VAR
In contrast to CAT-DIFF we then select data points around the median with a range according to the pruning percentage that we want to achieve, e.g. data points between the 25th and 75th percentile for a pruning percentage of 50%. This aligns with several works which have found variance in model signal to be predictive of difficulty (Swayamdipta et al., 2020; Agarwal et al., 2022a). Swayamdipta et al. (2020) show that data points that fall on the tails of the variance tend to be either too easy or too difficult and thus the model may not benefit most from being trained on such examples.
@@ 4.2 Evaluation Data
Our evaluation sets consist of WMT18 (Bojar et al., 2018), WMT15 Bojar et al. (2015), FLORES (Goyal et al., 2022; Costa-jussà et al., 2022) and MAFAND-MT (Adelani et al., 2022a) test sets. These test sets were curated using human annotators thereby mitigating common issues associated with automatically aligned datasets, such as erroneous alignments. The WMT18 and WMT15 test sets were released as part of the WMT machine translation task consisting of 2998 and 1500 parallel sentences respectively from various sources of the news domain. WMT18 was used to evaluate German models while WMT15 test set was used to evaluate French. The FLORES test set was used to evaluate all languages: German, French and Swahili. The FLORES test set consists of 1012 sentences from 3 sources: WikiNews, WikiJunior and WikiVoyage. We use MAFAND-MT test set for Swahili model evaluation. MAFAND-MT test set consists of 1835 sentences from the news domain.
@@ 4.3 Training Details
We train all our NMT models using the fairseq library [^2] https://github.com/facebookresearch/fairseq (Ott et al., 2019). Our models follow the same architecture as the vanilla transformer (Vaswani et al., 2017) with GeLU activation (Hendrycks & Gimpel, 2016) instead of ReLU activation (Nair & Hinton, 2010). We report all the hyperparameters used for training the models in Table 1. We conduct all our experiments on 4 A100 40GB GPUs with the duration for runs at 90%, 70%, 50% prune level, and full dataset utilization approximately taking 1, 2, 3, and 4.5 hours, respectively. Furthermore, we train separate SentencePiece tokenizers (Kudo & Richardson, 2018) for source and target languages with a vocab size of 16K using 3.8M translation pairs for both English-German and English-Swahili experiments. We use SacreBLEU library [^3] With parameters: nrefs:1|case:mixed|eff:no|tok:13a|smooth:exp|version:2.3.2 (Post, 2018) to compute BLEU score (Papineni et al., 2002) and used it as our evaluation metric. We also calculate chrF++ and COMET scores and report them in the Appendix. All the metrics exhibit the same trend in our experiments. Our baseline models are trained on the entire 3.8M translation pairs and denoted as “full” for all language pairs; English-German, English-French and English-Swahili. We investigate the impact of data pruning with sparsity levels of 50%, 70% and 90% using various techniques briefly described below.
@@ 4.4 Baselines
    Random Selection: This involves choosing data instances without specific criteria purely by chance. Although this approach may seem intuitively suboptimal, both Azeemi et al. (2023) and Marion et al. (2023) show that this also performs competitively with different pruning strategies for few cases. Random selection is also consistently reported in other problems evaluating the ability of a method to identify a salient subset such as weight pruning (Gale et al., 2019) or interpretability (Hooker et al., 2019). Therefore, we include this approach as a meaningful expected lower bound of performance for any pruning method proposed. This helps us answer the question, is this method better than randomly guessing what data to keep?
@@
    COMET-QE: Rei et al. (2020) proposed model-based MT evaluation metric that utilises embeddings from XLM-RoBERTa Conneau et al. (2020) to access translation quality when provided source, labels and target translations or just source and target translations.
@@
    LaBSE & LASER: Both are multilingual sentence embedding models trained on large aligned corpora covering a number of languages. LASER (Heffernan et al., 2022) was covers 200 languages and was trained with an encoder-decoder LSTM architecture. LaBSE (Feng et al., 2022) on the other hand covers 93 languages and was trained on top of BERT Devlin et al. (2019).
@@ 5 Results and Discussion
Figure 1: Performance of the CAT methods for both German and Swahili. The CAT-DIFF(1,5), referring to difference between checkpoint 1 and 5, consistently performs better than random. CAT-VAR yields very poor performance in Swahili and also underperforms CAT-DIFF for German.
@@ 5.1 CAT techniques outperforms random pruning in German and Swahili
Figure 1(a) demonstrates that CAT-DIFF(1,5) and CAT-VAR outperform random selection in the German test sets. On the other hand, in Swahili, only CAT-DIFF(1,5) consistently outperforms random selection, while selecting data randomly yields better results than CAT-VAR (refer to Figure 1(b)). CAT techniques achieve efficient pruning while maintaining an upwards of 75% performance compared to training on the full dataset, even after pruning 90% of the data. For German, both CAT-DIFF(1,5) and CAT-VAR maintain over 80% of the performance achieved by training on 100% of the data. In the case of Swahili, CAT-DIFF(1,5) retains above 80% of the performance, whereas CAT-VAR tends to perform worse, retaining only around 30% of the performance. A noteworthy efficiency of CAT techniques lies in their ability to sidestep the inferencing with fully converged models, thus one doesn’t need a lot of compute to utilise these techniques.
Table 2 presents the ablation study conducted for CAT-DIFF to compare the performance of CAT pruning techniques based on different checkpoints. We evaluate the difference between checkpoint 1 and 5, referred to as CAT-DIFF(1,5) and checkpoint 1 and 2, denoted as CAT-DIFF(1,2). We see that CAT-DIFF(1,5) yield better performance than CAT-DIFF(1,2). However, CAT-DIFF(1,2) offers the advantage of using fewer resources albeit at the cost of a slight dip in performance.

### v2

@@ Abstract
Neural Machine Translation models are extremely data and compute-hungry. However, not all data points contribute equally to model training and generalization. Data pruning to remove the low-value data points has the benefit of drastically reducing the compute budget without a significant drop in model performance. In this paper, we propose a new data pruning technique: Checkpoints Across Time (CAT), that leverages early model training dynamics to identify the most relevant data points for model performance. We benchmark CAT against several data pruning techniques including COMET-QE, LASER and LaBSE. We find that CAT outperforms the benchmarks on Indo-European languages on multiple test sets. When applied to English-German, English-French and English-Swahili translation tasks, CAT achieves comparable performance to using the full dataset, while pruning up to 50% of training data. We inspect the data points that CAT selects and find that it tends to favor longer sentences and sentences with unique or rare words.
@@ 1 Introduction
In recent years, there has been significant improvement in the quality of Neural Machine Translation (NMT) models (Johnson et al., 2017; Arivazhagan et al., 2019; Liu et al., 2020; Fan et al., 2021; Reid et al., 2021; Ramesh et al., 2022; Costa-jussà et al., 2022; Bapna et al., 2022; Gala et al., 2023). The success of these models is primarily due to many factors including pretraining data size, compute abundance and the ever-increasing model size. Despite the gains due to growing datasets, large dataset sizes have also posed significant hurdles for maintaining data quality. Significant portions of web-scraped data used for language model pretraining have been shown to be of low quality, machine-generated spam, pornographic content (Kreutzer et al., 2022). Given the ever-growing size of parallel corpora, it often becomes laborious for humans to assess the quality at scale (Freitag et al., 2021; Agarwal et al., 2022; Longpre et al., 2023). Quality of pretraining data for large NMT models often faces amplified quality issues because of extensive bitext mining approaches to automatically extracting translation pairs from either monolingual corpora (Guo et al., 2018; Schwenk et al., 2021a, b; Costa-jussà et al., 2022; Ramesh et al., 2022; Vegi et al., 2022; Gala et al., 2023) or document-aligned corpora (Bañón et al., 2020; Steingrímsson et al., 2021; Ramesh et al., 2022; Gala et al., 2023). Bitext mining approaches, although far more cost-effective than human curation, have been shown to introduce significant noise into datasets by relying on sub-optimal sentence embedding models (Thompson & Koehn, 2019; Feng et al., 2022; Heffernan et al., 2022) to match sub-optimal translation pairs (Kreutzer et al., 2022).
@@
CAT is motivated by fundamental work in machine learning which has shown that there are distinctive periods to learning in deep neural network training. Prior work shows that more common, easier features are learned earlier in training, with the most challenging features learned in the last stages of training (Jiang et al., 2020; Achille et al., 2017; Siddiqui et al., 2022). Prior work has also shown that variance in gradients with respect to inputs (Agarwal et al., 2022) or gradient patterns themselves vary depending on the period of learning (Faghri et al., 2020). However, gradients remain expensive to leverage as a learning signal to identify relevant instances. Furthermore, almost the entirety of this work has focused on a computer vision setting, with the exception of Swayamdipta et al. (2020) which uses differences in the model’s confidence in the true class, and the variability of this confidence across epoch to cluster data. Here – we explore applying the insights that there are distinctive periods of training to an NLP setting and leveraging for the purpose of pruning. We propose a more efficient method to compute signal relative to prior methods which have required computing gradients instead. Our hypothesis is that those with the largest variability indicate the easy examples which show the quickest learning and gains in certainty early in training.
@@
We study the effectiveness of CAT across diverse linguistic contexts. Specifically, we utilize English-German and English-French pairs from the Indo-European family using WMT datasets, and the English-Swahili pair from the Bantu family with automatically aligned datasets. Through a series of extensive experiments, we show that CAT results in significant improvements in translation quality over random pruning (randomly selecting $X\%$ subset of the training data) and other embedding-based quality estimation methods such as LaBSE (Feng et al., 2022). We summarize our key contributions below:
@@
    We propose a novel method, Checkpoints Across Time (CAT), that leverages perplexity variation to prune MT datasets with significant improvements over existing methods for Indo-European languages. We show that CAT results in significant improvements in translation quality over random pruning (randomly selecting $X\%$ subset of the training data) and other embedding-based quality estimation methods such as LaBSE (Feng et al., 2022). Using our best CAT-based method on English-German, English-French and English-Swahili translation tasks, we are able to achieve comparable performance while pruning up to 50% of the training data.
@@ 2.1 Data Pruning in ML
Data pruning has been extensively studied in the work of Marion et al. (2023); Li et al. (2018); Raju et al. (2021); Swayamdipta et al. (2020); Boubdir et al. (2023). Li et al. (2018); Sorscher et al. (2022) and Raju et al. (2021) focus on data pruning for computer vision whereas Swayamdipta et al. (2020) focuses on data pruning for language tasks. Various techniques have been proposed for data pruning, such as utilization of loss profile (Siddiqui et al., 2022), gradients confidence score (Agarwal et al., 2022), and data classification using a model (Swayamdipta et al., 2020). Increasingly techniques have been extended from primarily computer vision settings (Sorscher et al., 2022) to more recently expanding the treatment of data pruning language settings (Marion et al., 2023). These methods quantify the importance of data points to model generalization. Our work follows along these lines of identifying critical data points by leveraging the difference in perplexity between early checkpoints.
@@ 2.2 Model-based Data Pruning for NMT
All these approaches rely on the existence of a small trusted data set or trusted models to extrinsically define what “good data” looks like. Our approach, while also using model perplexity as a metric, however, does not require such pre-selection, and focuses on model training dynamics instead. Thereby, it can be applied more broadly. Our experiments aim at this less constrained setting without any given high-quality data, so we compare against model-based selection with an MT-QE model and a multilingual LLM.
@@ 2.4 Comparison Studies
With this plethora of filtering methods and application scenarios, there have been few works to systematically compare them. The WMT shared tasks for corpus filtering (Koehn et al., 2018, 2019, 2020) have shed light on how different languages and resourcefulness conditions pose different challenges for above-described filtering methods in practice, and how they affect downstream NMT performance. Bane & Zaretskaya (2021) compare data filtering techniques with a more qualitative angle: They find that NMT scores and MUSE embeddings have the highest correlation with human quality judgments, NMT scores work best for in-domain selection, and MUSE/XLM-R for out-of-domain generalization. Herold et al. (2022) compare multiple data noise detection techniques according to their effectiveness for filtering out specific types of data noise. They find that LASER does not detect incorrect languages, and both LASER and cross-entropy perform weakly on detecting misaligned sentences and over/under translation. Similarly, Bane et al. (2022) find cross-entropy filtering empirically superior to other methods (XLM-R, MUSE, LASER, COMET) for filtering out various synthetic noise types (Khayrallah & Koehn, 2018). In their setup, COMET fails at filtering out misaligned sentences but shows particular sensitivity for target-side omissions, NMT scoring for source-side omissions, while LASER and COMET do not filter mismatching numbers well. Dakwale et al. (2022) compare LASER and LaBSE on low-resource corpus filtering tasks and find them competitive if languages are included in the pretraining of the underlying embedding models.
@@ 3 CAT: Checkpoints Across Time
Our goal is to extract the most valuable data points that would contribute to model generalization and train the model solely on this. Our approach, while simple, leverages the difference in learning stages in deep neural network optimization. Firstly, we train an NMT model on the full dataset for only a few epochs, ideally a small fraction of the total time needed for the model to converge. Subsequently, we use the checkpoints from this initial stage to compute the perplexity of each data point. These perplexity scores serve as the basis for selecting candidate data points for subsequent pruning using different perplexity profiles. This approach draws upon the findings of Swayamdipta et al. (2020) that categorize training data relevance based on the learning dynamics during training, and Agarwal et al. (2022) which leverage the difference in gradients across training to identify easy versus challenging examples. Unlike Agarwal et al. (2022) which uses computationally expensive variance of gradients, we use perplexity. For CAT we can utilize the perplexity scores in two ways:
@@ 3.1 CAT-DIFF
From the ranked data points, we select those with the highest differences between early epochs, [^1] Either 5th and 1st or 2nd and 1st. according to a threshold at a chosen percentile. For example, for pruning of 50% of data points, we keep all data points that rank higher than the 50th percentile.
@@ 3.2 CAT-VAR
In contrast to CAT-DIFF we then select data points around the median with a range according to the pruning percentage that we want to achieve, e.g. data points between the 25th and 75th percentile for a pruning percentage of 50%. This aligns with several works which have found variance in model signal to be predictive of difficulty (Swayamdipta et al., 2020; Agarwal et al., 2022). Swayamdipta et al. (2020) show that data points that fall on the tails of the variance tend to be either too easy or too difficult and thus the model may not benefit most from being trained on such examples.
@@ 4.2 Evaluation Data
Our evaluation sets consist of WMT18 (Bojar et al., 2018), WMT15 Bojar et al. (2015), FLORES (Goyal et al., 2022; Costa-jussà et al., 2022) and MAFAND-MT (Adelani et al., 2022a) test sets. These test sets were curated using human annotators thereby mitigating common issues associated with automatically aligned datasets, such as erroneous alignments. The WMT18 and WMT15 test sets were released as part of the WMT machine translation task consisting of 2998 and 1500 parallel sentences respectively from various sources of the news domain. WMT18 was used to evaluate German models while WMT15 test set was used to evaluate French. The FLORES test set was used to evaluate all languages: German, French and Swahili. The FLORES test set consists of 1012 sentences from 3 sources: WikiNews, WikiJunior and WikiVoyage. We use the MAFAND-MT test set for Swahili model evaluation. MAFAND-MT test set consists of 1835 sentences from the news domain.
@@ 4.3 Training Details
We train all our NMT models using the fairseq library [^2] https://github.com/facebookresearch/fairseq (Ott et al., 2019). Our models follow the same architecture as the vanilla transformer (Vaswani et al., 2017) with GeLU activation (Hendrycks & Gimpel, 2016) instead of ReLU activation (Nair & Hinton, 2010). We report all the hyperparameters used for training the models in Table 1. We conduct all our experiments on 4 A100 40GB GPUs with the duration for runs at 90%, 70%, and 50% prune level, and full dataset utilization approximately taking 1, 2, 3, and 4.5 hours, respectively. Furthermore, we train separate SentencePiece tokenizers (Kudo & Richardson, 2018) for source and target languages with a vocab size of 16K using 3.8M translation pairs for both English-German and English-Swahili experiments. We use SacreBLEU library [^3] With parameters: nrefs:1|case:mixed|eff:no|tok:13a|smooth:exp|version:2.3.2 (Post, 2018) to compute BLEU score (Papineni et al., 2002) and used it as our evaluation metric. We also calculate chrF++ and COMET scores and report them in the Appendix. All the metrics exhibit the same trend in our experiments. Our baseline models are trained on the entire 3.8M translation pairs and denoted as “full” for all language pairs; English-German, English-French and English-Swahili. We investigate the impact of data pruning with sparsity levels of 50%, 70% and 90% using various techniques briefly described below.
@@ 4.4 Baselines
    Random Selection: This involves choosing data instances without specific criteria purely by chance. Although this approach may seem intuitively suboptimal, both Azeemi et al. (2023) and Marion et al. (2023) show that this also performs competitively with different pruning strategies for a few cases. Random selection is also consistently reported in other problems evaluating the ability of a method to identify a salient subset such as weight pruning (Gale et al., 2019) or interpretability (Hooker et al., 2019). Therefore, we include this approach as a meaningful expected lower bound of performance for any pruning method proposed. This helps us answer the question, is this method better than randomly guessing what data to keep?
@@
    COMET-QE: Rei et al. (2020) proposed model-based MT evaluation metric that utilizes embeddings from XLM-RoBERTa Conneau et al. (2020) to access translation quality when provided source, labels and target translations or just source and target translations.
@@
    LaBSE & LASER: Both are multilingual sentence embedding models trained on large aligned corpora covering a number of languages. LASER (Heffernan et al., 2022) covers 200 languages and was trained with an encoder-decoder LSTM architecture. LaBSE (Feng et al., 2022) on the other hand covers 93 languages and was trained on top of BERT Devlin et al. (2019).
@@ 5 Results and Discussion
Figure 1: Performance of the CAT methods for both German and Swahili. The CAT-DIFF(1,5), referring to the difference between checkpoints 1 and 5, consistently performs better than random. CAT-VAR yields very poor performance for Swahili and also underperforms CAT-DIFF for German.
@@ 5.1 CAT techniques outperforms random pruning in German and Swahili
Figure 1(a) demonstrates that CAT-DIFF(1,5) and CAT-VAR outperform random selection in the German test sets. On the other hand, in Swahili, only CAT-DIFF(1,5) consistently outperforms random selection, while selecting data randomly yields better results than CAT-VAR (refer to Figure 1(b)). CAT techniques achieve efficient pruning while maintaining an upwards of 75% performance compared to training on the full dataset, even after pruning 90% of the data. For German, both CAT-DIFF(1,5) and CAT-VAR maintain over 80% of the performance achieved by training on 100% of the data. In the case of Swahili, CAT-DIFF(1,5) retains above 80% of the performance, whereas CAT-VAR tends to perform worse, retaining only around 30% of the performance. A noteworthy efficiency of CAT techniques lies in their ability to sidestep the inferencing with fully converged models, thus one doesn’t need a lot of compute to utilize these techniques.
Table 2 presents the ablation study conducted for CAT-DIFF to compare the performance of CAT pruning techniques based on different checkpoints. We evaluate the difference between checkpoints 1 and 5, referred to as CAT-DIFF(1,5) and checkpoints 1 and 2, denoted as CAT-DIFF(1,2). We see that CAT-DIFF(1,5) yields better performance than CAT-DIFF(1,2). However, CAT-DIFF(1,2) offers the advantage of using fewer resources albeit at the cost of a slight dip in performance.